In the Bohr model of the atom, an electron can be thought of a small sphere that rotates around the nucleus. In a hydrogen atom, an electron (me=9.11 x 10^-31 kg) orbits a proton at a distance of 5.3 x 10^-11 m from the proton. If the proton pulls on the electron with a force of 9.2 x 10^8 N, how many revolutions per second does the electron make?

6. In the Bohr model of the atom, an electron can be thought of a small sphere that rotates around the nucleus. In a hydrogen atom, an electron (me-9.11 × 10-31 kg) orbits a proton at around the nucleus. In a hydrogen atom, an electron (me 9.11 x 10-31 kg) orbits a proton at distance of 5.3 × 10-11 rn from the proton. If the proton pulls on the el 9.2 x 108 N, how many revolutions per second does the electron make? ectron with a force of
Answer #1

m_e=9.11\times 10^{-31}kg

d=5.3\times 10^{-11}m

F=9.2\times 10^8 N

This problem could be treated like a circular motion, the attracting force due to the electromagnetism is the centripetal force.

The centripetal force is related to the angular velocity through the formula

F_c=m.\omega ^2.R

as:

m= mass

w= angular velocity

R= radius

We need to now the angular velocity (how many revolutions per second does the electron make)

clearing w

\omega ^2=\frac{F_c}{m.R}

\omega =\sqrt{\frac{F_c}{m.R}}

\omega =\sqrt{\frac{9.2\times10^8N}{9.11\times10^{-31}Kg.5.3\times10^{-11}m}}

\omega =4.365\times10^{24}rev/s

/*******************************************************************

If you have any question about it please leave me a comment

> It's the correct answer, but you forget to convert the angular velocity from rad/s to rev/s by dividing by 2π in order to get 6.947 x 10^23 rev/s

TonySuprano Tue, Mar 1, 2022 12:13 PM